Question 9 of 12 Submit What is the pH of a 9.7 x 10-6 M KOH solution? 1 4 7 +/- 2 5 8 : 3 6 9 0 x 10
Question 10 of 12 Submit The pOH of 0.0240 M solution of Sr(OH)2 is 1 4 7 +/- 2 5 8 : 3 6 9 0 x 10
Answer #1

O = 9.7x106M Concentration g kom - [DOH = -log [HⓇ] = -log[9.7x106]. 6-log 9.7 . : = 6-0.9867 = 5.01 .. pH = 14 -POH = 14 - 5